# fastrerandomize: An R Package for Fast Rerandomization Using Accelerated Computing

### v1

@@ 3.4 Hardware Acceleration
Figure 5: Comparing hardware-accelerated performance of rerandomization steps on the CPU versus GPU. Bars show elapsed times (seconds) for (top) generating large pools of randomizations and (bottom) conducting the randomization-based inference. Left/right varies the maximum number of Monte Carlo draws (either $1\times 10^{5}$ or $2\times 10^{5}$). The $x$-axis indicates the number of covariates, the $y$-axis runtime. GPU acceleration provides substantial speed-ups in both the generation and testing phases, particularly as the number of covariates or draws increases. Here, $n$=1000; the CPU backend is an Apple M series chip (14 core); the GPU is 40-core METAL backend device.
Figure 5 illustrates how GPU acceleration yields substantial performance gains relative to the CPU for both randomization generation and randomization testing. With a larger number of covariates ($1000$) or more intensive Monte Carlo sampling ($2\times 10^{5}$ draws), GPU acceleration reduces the elapsed time dramatically by executing linear algebra and sampling tasks across thousands of cores in parallel. This allows the software to handle large-scale rerandomization designs that would otherwise be computationally prohibitive under a CPU-only workflow. (Note, though, that CPU-only users can still benefit from JIT-compilation and automatic vectorization capabilities in the package.)
Next, Figure 6 illustrates how the GPU-vs.-CPU speedup (i.e., how many times faster GPU is relative to CPU) increases when moving from $n=100$ to $n=1000$. On the horizontal axis, we vary the number of covariates; on the vertical axis, we plot the ratio of GPU speedups:
$\frac{\bigl{(}\text{CPU time}/\text{GPU time}\bigr{)}_{\!n=1000}}{\bigl{(}\text{CPU time}/\text{GPU time}\bigr{)}_{\!n=100}}.$
Values above 1 indicate the GPU’s performance advantage becomes larger at $n=1000$ than at $n=100$. The figure is faceted by the two main package steps — generate_randomizations vs. randomization_test—and by the number of Monte Carlo draws ($1\times 10^{5}$ or $2\times 10^{5}$). As the problem size grows, the ratio typically exceeds 1 (particularly so for the more computationally expensive candidate randomization generation step), indicating that the GPU’s relative speedup improves with larger $n$.
@@
4 FastRerandomize Package Use
fastrerandomize provides an interface for researchers to (1) install and configure a hardware-accelerated JAX backend, (2) generate pools of acceptable randomizations via exact enumeration or Monte Carlo sampling, and (3) perform randomization-based inference respecting rerandomization. This section outlines the basic steps to install and initialize the software, generate randomizations, and conduct inference, followed by advanced usage tips for large-scale designs and pre-analysis evaluation.
@@ 4.1 Package Installation and Loading
The package automatically calls out to JAX for hardware-accelerated computations. By default, fastrerandomize attempts to locate an existing conda environment named "fastrerandomize". To create or update this environment with all necessary backend packages, run:
@@
This function installs numpy, jax, and other dependencies in a minimal Python 3.11 environment via conda. If you have a CUDA-capable GPU and drivers properly configured, jaxlib with CUDA support can be employed. On Mac M+ architectures, jax-metal is installed for METAL-based GPU acceleration.
@@ 4.2 Key Functions Overview
Creates or updates a conda environment for running the hardware-accelerated JAX backend. By default, the environment is named "fastrerandomize". Internally, this function installs (or upgrades) jax, jaxlib, and related dependencies in a lightweight Python 3.11 environment. On machines with NVIDIA GPUs, a CUDA-enabled jaxlib is installed automatically if CUDA drivers are detected; on Apple Silicon (M1/M2), the METAL-accelerated jax is installed instead. This function is typically run once per session before calling any other fastrerandomize routines. The key arguments include:
@@
  • force: Logical. If TRUE, forces an environment rebuild even if one already exists.
@@
Produces a pool of acceptable randomizations (treatment assignments) under a specified experimental design, employing either exact enumeration or Monte Carlo sampling. Internally, the function leverages JAX-based just-in-time (JIT) compilation to accelerate repeated operations (like random sampling and covariate-balance checks) on CPU or GPU. Accepted designs are retained based on a user-specified rerandomization threshold (e.g., Hotelling’s $T^{2}$ or a custom balance statistic). The output is returned as an S3 object of class fastrerandomize_randomizations, which has its own summary() and plot() methods.
@@
Together, these functions support an end-to-end workflow for rerandomization:
@@
    Building a JAX-accelerated backend with build_backend();
@@
Below, we illustrate these usage patterns in practice before later presenting two substantive applications involving real RCT data.
4.3 Workflow
1. Building the JAX backend.
@@ 4.3 Workflow
When first working with the fastrerandomize packages, users should construct the environment and install JAX using build_backend():
@@
Figure 8: Result of plot(test_results).
4.4 Advanced Usage and Practical Considerations
@@
Selecting an acceptance threshold,
  • randomization_accept_prob
balances tighter covariate balance (lower acceptance probability) against the size of the retained randomization pool. If the threshold is too restrictive, you might have too few randomizations for stable inference. Conversely, if it is too lenient, the statistical benefits of rerandomization are diminished. A recommended approach is to explore multiple acceptance probabilities and compare how each influences downstream inference.
@@
You can then load only subsets of the file into R for downstream analysis, preserving memory. Similarly, the return_type parameter can be set to "jax" if your workflow stays mostly in JAX (advanced usage).
@@
By default, the package uses Hotelling’s $T^{2}$ as the rerandomization criterion. You can supply a different function to threshold_func in generate_randomizations() if you prefer alternative balance statistics. The function must be vectorized in JAX (via vmap), returning a numeric measure for each randomization.
@@
When findFI = TRUE in randomization_test(), the package attempts to compute an exact rerandomization-based interval. For large sample sizes or extremely large randomization pools, this calculation can be time-consuming. It is often helpful to set a smaller acceptance threshold or restrict the size of the randomization set when first experimenting with fiducial interval estimation.
4.5 Advanced Use: Optimal Threshold Exploration
@@
4.6 Summary of Workflow
Figure 1 and the examples above illustrate the recommended workflow:
  1. 1.
    Initialize and build the Python/JAX backend, enabling JIT compilation and optional TPU/GPU acceleration.
  2. 2.
    Simulate or load covariate data for the planned experiment.
  3. 3.
    Generate randomizations by exact enumeration (small $n$) or Monte Carlo sampling (large $n$), using compiled code and, if available, GPU-accelerated batching.
  4. 4.
    Collect the resulting set of acceptable assignments, which can be returned to R or saved to disk (again exploiting compiled JAX routines).
  5. 5.
    Perform randomization-based inference using randomization_test() for your observed outcomes, respecting the rerandomization design and reusing the same GPU-accelerated kernels.
By following these steps, researchers can incorporate rerandomization into their experimental workflow while harnessing hardware-accelerated computations to handle designs that would otherwise be prohibitively expensive computationally.
5 Examples
5.1 Substantive Example 1: An Agricultural Field Experiment
In what follows, we illustrate a complete workflow for an agricultural field experiment drawn from
@@ 5.1 Substantive Example 1: An Agricultural Field Experiment
# Setup: install, load package, initialize JAX
@@
# install.packages("devtools")
# devtools::install_github("cjerzak/fastrerandomize-software/fastrerandomize")
@@
# Optional: build the Python/JAX backend in a conda environment
@@
candidate_randomizations = rand_mc\$randomizations,
@@
5.2 Substantive Example 2: An Anti-Poverty RCT
To enrich the original anti-poverty dataset, we followed the approach of Jerzak et al. (2023) and appended pre-treatment image features derived from satellite imagery centered on each RCT unit’s approximate geo-location. This augmentation yields over 500 unique covariates per unit from a fine-tuned CLIP image model (Lu et al., 2017), capturing a fine-grained depiction of pre-treatment contextual conditions (e.g., related to the environment or agriculture). We then assess the randomness of various treatment assignments in this high-dimensional space by computing the Mahalanobis distance, which serves as a robust measure of covariate balance across treatment and control groups.
The below codeblock will compare Mahalanobis distance values from the original experiment with what we would obtain via rerandomization in image space. A greater Mahalanobis distance value is an indicator that the covariate means for the treatment and control groups are very different, meaning that the randomization did not produce groups with well-balanced covariates. Thus, “good” randomizations ought to minimize their M-distance values. We will compare the M-distance value under the implemented and a hypothetical rerandomization-based design.
# Acquire package data for the Youth Opportunities Experiment
data( YOPData, package = "fastrerandomize")
#Make sure row dimensions match
dim(YOPData\$ImageEmbeddings)
dim(YOPData\$RCTData)
library(fastrerandomize)
set.seed(543210)
# Load the built-in data, which has:
# 1) YOPData\$RCTData (contains treatment indicator + possibly other columns)
# 2) YOPData\$ImageEmbeddings (CLIP-based satellite imagery embeddings)
data(YOPData)
# Verify the dimensions
dim(YOPData\$ImageEmbeddings) # e.g., N x K matrix of embeddings
dim(YOPData\$RCTData) # e.g., N x a few columns (including "treated")
# 1) Define a function to compute Mahalanobis distance
# between average treated and control covariates
compute_m_dist <- function(merged_assignment) {
library(dplyr)
# merged_assignment is a data.frame or matrix whose first column is "treated"
# and the remaining columns are the covariates/embeddings.
# The function then computes the sample mean difference,
# the pooled covariance matrix, and uses mahalanobis() to get the M-dist.
treated <- merged_assignment %>filter(treated == 1) %>select(-treated)
control <- merged_assignment %>filter(treated == 0) %>select(-treated)
m_distance <- mahalanobis(center = colMeans(treated),
x = colMeans(control),
cov = diag(diag(cov(rbind(treated, control)))))
return(m_distance)
# 2) Generate a rerandomization-based design:
# We’ll accept random assignments to T=1 or T=0
# only if they pass a certain (low) acceptance probability.
# This is just a demonstration: randomization_accept_prob = 0.0001
candidate_randomizations_array <- generate_randomizations(
n_units = length(YOPData\$RCTData\$treated),
n_treated = sum(YOPData\$RCTData\$treated),
X = YOPData\$ImageEmbeddings,
randomization_accept_prob = 0.0001, # strict acceptance
max_draws = 10^5,
batch_size = 10^4,
approximate_inv = TRUE
# This returns an S3 object of class "fastrerandomize_randomizations",
# which has (1) randomizations, (2) balance, etc.
cat("Number of accepted randomizations:",
[…]
# Compare M-distances
# (A) M-distance of a purely random assignment
# 1) random draw of 0/1
# 2) cbind to embeddings
# 3) compute M-distance
random_array <- sample(YOPData\$RCTData\$treated,
size = nrow(YOPData\$ImageEmbeddings),
replace = FALSE)
merged_arr <- cbind("treated" = random_array,
YOPData\$ImageEmbeddings)
# (B) M-distance of the YOP Data’s actual assignment
# We assume the column in RCTData is called "treated" (if not, rename below).
merged_assignment <- cbind("treated" = YOPData\$RCTData\$treated,
YOPData\$ImageEmbeddings)
# (C) Average M-distance across the accepted randomizations
# We’ll loop over some of the accepted
# assignment vector in candidate_randomizations_array
# and compute the M-dist. Then we compute the average
nM <- 100 # how many accepted to analyze
SampIndices <- sample(1:nrow(candidate_randomizations_array\$randomizations),
nM,
replace = T)
mean_m_dist <- 0
mDist_vec <- c()
for (i in SampIndices){
# For each row i, we have a length-N vector of 0/1 for the assignment
merged_temp <- as.data.frame(
YOPData\$ImageEmbeddings))
mDist_ <- compute_m_dist(merged_temp)
mDist_vec <- c(mDist_vec, mDist_)
mean_m_dist <- mean_m_dist + mDist_ / nM
# print results
[…]
# Random assignment M-distance: 1.506205
[…]
# Actual YOPData assignment M-distance: 13.5678
[…]
# M-distance over accepted randomizations: 0.3384692
In terms of results, we see from these M-distance calculations that the purely random assignment yields a moderate degree of imbalance in the high-dimensional covariate space, while the actual YOP assignment exhibits an even larger imbalance (M-distance of 13.57). In contrast, the rerandomized assignments accepted by fastrerandomize achieve substantially lower M-distances (mean of 0.34 among the sampled accepted assignments).
We also found that the calculation of just $10^{2}$ randomization M distances in this high-dimensional case in (admittedly un-optimized) base R has a longer runtime than the hardware accelerated calculation of $10^{5}$ distances using fastrerandomize. These results highlight how hardware-accelerated rerandomization can markedly improve covariate balance when dealing with high-dimensional embeddings (e.g., hundreds of satellite-derived features). Simply put, the package can help unlock more stringent rerandomization designs for more complex, high-dimensional settings.
# Performing Randomization Tests
random_test_one <- fastrerandomize::randomization_test(
X = YOPData\$ImageEmbeddings,
obsY = YOPData\$RCTData[,"outcome"],
obsW = candidate_randomizations_array\$randomizations[1,],
random_test_one\$p_value # 0.2
summary(random_test_one)
6 Future Development Plans
Although the current iteration of fastrerandomize provides a robust and scalable tool for conducting rerandomization in large-scale experiments, several avenues for future development remain. Our primary goals include:
7 Conclusion
The fastrerandomize package provides fast rerandomization and randomization inference by combining a statistically principled framework with hardware-accelerated computations. While the current release already supports large-scale designs by leveraging efficient memory operations and JAX backends (CPU, CUDA, and METAL), future work will focus on scaling to even larger randomization spaces by refining memory handling, exploring approximate methods that reduce the computational footprint for massive experiments, accommodating additional balance criteria to reflect more nuanced research needs, and broadening hardware acceleration capabilities across new platforms. By pursuing these directions, we aim to solidify fastrerandomize as a robust solution for researchers facing demanding experimental designs that require rigorous randomization-based inference.
Computational Details
License
Acknowledgments
Authors listed in alphabetical order. We thank Kaz Sakamoto and the R community for helpful feedback and testing.

### v2

@@ 4.2 Simulation Study Results
Figure 6: Comparing performance of rerandomization steps on the CPU versus GPU for $n=100$. Bars show elapsed times (seconds) for (top) generating large pools of randomizations and (bottom) conducting the randomization-based inference. Left/right varies the maximum number of Monte Carlo draws (either $1\times 10^{5}$ or $2\times 10^{5}$). The $x$-axis indicates the number of covariates, the $y$-axis runtime. GPU acceleration provides substantial speed-ups in both the generation and testing phases, particularly as the number of covariates or draws increases. The CPU backend is an Apple M series chip (14 core); the GPU is 40-core METAL backend device.
Next, Figure 7 depicts analogous performance measures for the larger sample size case ($n=1000$). Recall that, while at $n=100$, it was still feasible (if slow) to apply the base R approach for higher-dimensional or larger-draw settings, at $n=1000$ this implementation becomes prohibitively expensive. fastrerandomize-GPU continues to scale well. For instance, with 1000 covariates and $2\times 10^{5}$ Monte Carlo draws, fastrerandomize-CPU requires over a minute to generate pools of acceptable assignments, but merely requires 5 seconds with GPU by exploiting massively parallel balance checks. This indicates how the benefits of hardware acceleration are greatest for the large $n$ and $k$ regimes in fastrerandomize.
Figure 7: Comparing performance of rerandomization steps on the CPU versus GPU for $n=1000$. Bars again show elapsed times (seconds) for (top) generating large pools of randomizations and (bottom) conducting the randomization-based inference. Left/right varies the maximum number of Monte Carlo draws (either $1\times 10^{5}$ or $2\times 10^{5}$). The $x$-axis indicates the number of covariates, the $y$-axis runtime.
Figure 8 further compares how GPU vs. CPU speedups themselves change when moving from $n=100$ to $n=1000$. For each combination of covariate dimension, draw count, and rerandomization step (pool generation or inference), we compute the CPU/GPU ratio at $n=100$ and at $n=1000$, and then form the ratio of these two speedups. Values above 1 indicate that the GPU advantage becomes more pronounced in larger-scale simulations, an effect that is most evident in generation of randomization pools. As the dimension and Monte Carlo draws grow, the parallelized linear algebra driving fastrerandomize-GPU increasingly dominates the CPU-based routines, producing faster execution and enabling the scale up of rerandomization in experiments in high dimensional regimes.
@@
Overall, these simulation studies confirm that fastrerandomize can substantially alleviate the computational burden of rerandomization and randomization-based inference, even under stringent acceptance thresholds or high covariate dimensions. Although the CPU-only backend already offers benefits via JIT compilation, the GPU backend particularly stands out when the number of units or covariates becomes large and the total space of candidate randomizations expands accordingly.
@@
5 FastRerandomize Use
fastrerandomize provides an interface for researchers to (1) install and configure backends for hardware acceleration, (2) generate pools of acceptable randomizations via exact enumeration or Monte Carlo sampling, and (3) perform randomization-based inference respecting rerandomization. This section outlines the basic steps to install and initialize the software, generate randomizations, and conduct inference, followed by advanced usage tips for large-scale designs and pre-analysis evaluation.
@@ 5.1 Package Installation and Loading
By default, fastrerandomize attempts to locate an existing conda environment named "fastrerandomize". To create or update this environment with all necessary backend packages, run:
@@
This function installs the backend. [^3] The backend includes numpy, jax, and other dependencies in a minimal Python 3.11 environment via conda. If you have a CUDA-capable GPU and drivers properly configured, jaxlib with CUDA support can be employed. On Mac M+ architectures, jax-metal is installed for METAL-based GPU acceleration.
@@ 5.2 Key Functions Overview
Creates or updates a conda environment for running the (potentially) hardware-accelerated backend. By default, the environment is named "fastrerandomize". Internally, this function installs (or upgrades) dependencies in a lightweight Python 3.11 environment that automatically detects if CUDA drivers or METAL-accelerated GPUs are available. This function is typically run once per session before calling any other fastrerandomize routines. The key arguments include:
@@
Produces a pool of acceptable randomizations (treatment assignments) under a specified experimental design, employing either exact enumeration or Monte Carlo sampling. Internally, the function leverages just-in-time (JIT) compilation to accelerate repeated operations (like random sampling and covariate-balance checks) on CPU or GPU. Accepted designs are retained based on a user-specified rerandomization threshold (e.g., Hotelling’s $T^{2}$ or a custom balance statistic). The output is returned as an S3 object of class fastrerandomize_randomizations, which has its own summary() and plot() methods.
@@
Together, these functions support this end-to-end workflow for rerandomization (illustrated in Figure 1):
@@
    Building the computational backend with build_backend();
@@
By following these steps, researchers can incorporate rerandomization into their experimental workflow while harnessing hardware-accelerated computations to handle designs that would otherwise be prohibitively expensive computationally. See Appendix B for more workflow information and Appendix C for advanced use considerations.
@@
Next, we illustrate these usage patterns in practice before later presenting a substantive application involving real RCT data.
6 Tutorial: An Anti-Poverty RCT
To enrich the original anti-poverty dataset, we followed the approach of Jerzak et al. (2023) and appended pre-treatment image features derived from satellite imagery centered on each RCT unit’s approximate geo-location. This augmentation yields over 500 unique covariates per unit from a fine-tuned CLIP image model (Lu et al., 2017), capturing a fine-grained depiction of pre-treatment contextual conditions (e.g., related to the environment and local agriculture patterns). We then assess the degree of imbalance in this high-dimensional space by computing the Mahalanobis distance, which serves as a measure of covariate balance across treatment and control groups.
We first set up the environment, ensuring the fastrerandomize package is available, and reproducibility is maintained through seed-setting. The built-in dataset, YOPData, comprises two components: randomized control trial (RCT) indicators and satellite-derived image embeddings capturing high-dimensional covariate spaces.
library(fastrerandomize)
# Acquire package data for the Youth Opportunities Experiment
data( YOPData, package = "fastrerandomize")
dim(YOPData\$ImageEmbeddings) # Verify embeddings dimension
dim(YOPData\$RCTData) # Verify RCT data dimension
We next employ a stringent rerandomization strategy to produce well-balanced treatment-control groups. Candidate randomizations are evaluated based on their covariate balance, retaining only those surpassing a strict acceptance threshold. This approach aims at substantial improvements over naive random assignment, particularly valuable in high-dimensional covariate scenarios.
candidate_randomizations_array <- generate_randomizations(
n_units = length(YOPData\$RCTData\$treated),
n_treated = sum(YOPData\$RCTData\$treated),
X = YOPData\$ImageEmbeddings,
randomization_accept_prob = 0.0001, # strict acceptance
max_draws = 10^5,
batch_size = 10^4,
approximate_inv = TRUE)
cat("Number of accepted randomizations:",
nrow(candidate_randomizations_array\$randomizations))
We proceed to evaluate and compare the quality of the rerandomization outcomes through computed Mahalanobis distances. Three distinct scenarios are assessed: a purely random assignment, the actual assignment from the YOPData, and assignments generated through the rerandomization approach.
# (A) M-distance of a purely random assignment
random_array <- sample(YOPData\$RCTData\$treated,
size = nrow(YOPData\$ImageEmbeddings),
replace = FALSE)
merged_arr <- cbind("treated" = random_array,
YOPData\$ImageEmbeddings)
# (B) M-distance of the YOP Data’s actual assignment
merged_assignment <- cbind("treated" = YOPData\$RCTData\$treated,
YOPData\$ImageEmbeddings)
# (C) Average M-distance across the accepted rerandomizations
nM <- 100 # how many accepted to analyze
SampIndices <- sample(1:nrow(candidate_randomizations_array\$randomizations), nM, replace = T)
Further analysis includes calculating the average Mahalanobis distance across multiple accepted randomizations. By examining these averages, we underscore the potential improvements achievable through rerandomization in reducing covariate imbalance.
mean_m_dist <- 0
mDist_vec <- c()
for (i in SampIndices){
# For each row i, we have a length-N vector of 0/1 for the assignment
merged_temp <- as.data.frame(
YOPData\$ImageEmbeddings))
mDist_ <- compute_m_dist(merged_temp)
mDist_vec <- c(mDist_vec, mDist_)
mean_m_dist <- mean_m_dist + mDist_ / nM
# print results
cat("New random assignment M-distance:", random_m_distance)
# Random assignment M-distance: 1.51
cat("Actual YOPData assignment M-distance:", given_m_distance)
# Actual YOPData assignment M-distance: 13.57
cat("M-distance over accepted randomizations:", mean_m_dist)
# M-distance over accepted randomizations: 0.34
In terms of results, we see from these M-distance calculations that the purely random assignment yields a moderate degree of imbalance in the high-dimensional covariate space, while the actual YOP assignment exhibits an even larger imbalance (M-distance of 13.57). In contrast, the rerandomized assignments accepted by fastrerandomize achieve substantially lower M-distances (mean of 0.34 among the sampled accepted assignments). Thus, without rerandomization, imbalance is marked; with rerandomization, imbalance is minimal, and fastrerandomzie can help unlock rerandomization designs in high-dimensional settings such as we have here.
In terms of compute time, we found that the calculation of just $10^{2}$ randomization M distances in this high-dimensional case in base R has a longer runtime than the accelerated calculation of $10^{5}$ distances using fastrerandomize. These results highlight how accelerated rerandomization can markedly improve covariate balance when dealing with high-dimensional embeddings (e.g., hundreds of satellite-derived features). In this way, the package can help unlock more stringent rerandomization designs in complex, high-dimensional settings.
# Performing Randomization Tests
random_test_one <- fastrerandomize::randomization_test(
X = YOPData\$ImageEmbeddings,
obsY = YOPData\$RCTData\$outcome,
obsW = candidate_randomizations_array\$randomizations[1,],
random_test_one\$p_value # 0.2
summary(random_test_one)
7 Future Development
Although the current iteration of fastrerandomize provides a robust and scalable tool for conducting rerandomization in large-scale experiments, several avenues for future development remain.
8 Conclusion
The fastrerandomize package provides fast rerandomization by combining a statistically principled framework with hardware-accelerated computations. While the current release already supports large-scale designs by leveraging compiler-level optimizations, efficient memory operations, and (optionally) parallel processing hardware backends (GPU/TPU), future work will focus on scaling to even larger randomization spaces by refining memory handling, exploring approximate methods that reduce the computational footprint for massive experiments, accommodating additional balance criteria to reflect more nuanced research needs, and broadening hardware acceleration capabilities across new platforms. By pursuing these directions, we aim to solidify fastrerandomize as a robust solution for researchers facing demanding experimental designs that require rigorous randomization-based inference.
Computational Details
License
Acknowledgments
Authors listed in alphabetical order. We thank Kaz Sakamoto for helpful feedback and testing. For more information, see fastrerandomize.github.io.
Appendix A CPU vs. GPU Analysis for Rerandomization
Suppose we wish to generate $M$ total randomizations, divided into batches of size $B$. Let $k=\alpha\,d$ be the per-draw cost of checking covariate balance, where $\alpha$ is the cost per covariate per draw and $d$ is the number of covariates. Because each batch processes $B$ candidate randomizations in one pass, the total CPU runtime can be approximated by
$T_{\mathrm{CPU}}\;\approx\;r_{\mathrm{CPU}}\;+\;\left(\frac{M}{B}\right)\bigl{(}k\,B\bigr{)}\;=\;r_{\mathrm{CPU}}\;+\;k\,M,$
where $r_{\mathrm{CPU}}$ denotes CPU overhead (e.g., scheduling or memory operations). By contrast, a GPU-based approach adds a separate overhead $r_{\mathrm{GPU}}$ (i.e., based on kernel launches, synchronization steps, data transfers) but greatly reduces the time to check each batch by an effective parallelization factor $\beta$. Hence,
$T_{\mathrm{GPU}}\;\approx\;r_{\mathrm{GPU}}\;+\;\left(\frac{M}{B}\right)\frac{k\,B}{\beta}\;=\;r_{\mathrm{GPU}}\;+\;\frac{k\,M}{\beta}.$
Because many GPU cores operate in parallel, the large term $k\,M$ shrinks by approximately $\beta$. As a result, the GPU speedup,
$\frac{T_{\mathrm{CPU}}}{T_{\mathrm{GPU}}}\;=\;\frac{r_{\mathrm{CPU}}+k\,M}{r_{\mathrm{GPU}}+\frac{k\,M}{\beta}},$
converges to $\beta$ when $k\,M$ dominates both $r_{\mathrm{CPU}}$ and $r_{\mathrm{GPU}}$. In other words, as the experiment grows either in the number of Monte Carlo draws $M$ or in the dimension $d$, batching these balance checks on a GPU increasingly outperforms the CPU, mitigating the otherwise large runtime explosion from the $k\,M$ term. Consequently, GPU-accelerated rerandomization remains a practical solution for fast generation and assessment of millions of candidate assignments, even when covariate dimension or sample size is large.
Appendix B FastRerandomize Workflow
1. Building the backend.
@@ Appendix B FastRerandomize Workflow
When first working with the fastrerandomize packages, users should construct the environment using build_backend():
@@
Appendix C Advanced Usage and Practical Considerations
@@
Selecting an acceptance threshold, randomization_accept_prob, balances tighter covariate balance (lower acceptance probability) against the size of the retained randomization pool. If the threshold is too restrictive, you might have too few randomizations for stable inference. Conversely, if it is too lenient, the statistical benefits of rerandomization are diminished. A recommended approach is to explore multiple acceptance probabilities and compare how each influences downstream inference.
@@
You can then load only subsets of the file into R for downstream analysis, preserving memory.
@@
By default, the package uses Hotelling’s $T^{2}$ as the rerandomization criterion. You can supply a different function to threshold_func in generate_randomizations() if you prefer alternative balance statistics. The function must be vectorized (via JAX’s vmap), returning a numeric measure for each randomization.
@@
When findFI = TRUE in randomization_test(), the package attempts to compute an exact rerandomization-based interval. For large sample sizes or extremely large randomization pools, this calculation can be time-consuming. It is often helpful to set a smaller acceptance threshold or restrict the size of the randomization set when first experimenting with fiducial interval estimation. See Appendix D for discussion of how to set optimal rerandomization thresholds.
Appendix D Advanced Use: Optimal Threshold Exploration
@@
Appendix E Additional Tutorial: An Agricultural Field Experiment
In this additional example, we illustrate a complete workflow for an agricultural field experiment drawn from
@@ Appendix E Additional Tutorial: An Agricultural Field Experiment
# Setup: install, load package, initialize the computational backend
@@
# install.packages("fastrerandomize")
@@
# Optional: build the backend in a conda environment
@@
candidate_randomizations = rand_mc\\$randomizations,
@@
Appendix F Interactive Demonstration on Hugging Face
An interactive demonstration of fastrerandomize is available at fastrerandomize.github.io/space, where users can upload or simulate high-dimensional covariate data, specify rerandomization parameters such as treatment shares and acceptance thresholds, and then evaluate resulting randomizations through intuitive visual outputs.
Internally, the tool employs a Shiny application linked to our hardware-accelerated backend, allowing the user to experience how fastrerandomize can efficiently generate well-balanced treatment assignments for moderate- or large-scale experiments.
Figure 10 shows a screenshot of the interface, highlighting how the user can toggle between Monte Carlo or exact enumeration, set a maximum number of randomization draws or batch sizes, and observe the resulting distributions of Hotelling’s $T^{2}$.
In addition, one can upload or simulate outcome data to conduct a randomization-based hypothesis test, including the option of computing fiducial intervals. This workflow offers a hands-on illustration of the performance advantages and correct design-based inferences enabled by fastrerandomize, while requiring no additional configuration beyond a web browser.
Figure 10: Web-based demonstration of fastrerandomize.